# Thermodynamic Phase Transitions and Quantum Entropy Corrections in the Simpson–Visser Regular Black Hole

### v1

@@ I Introduction
The unresolved tension between general relativity and quantum mechanics stands as the central challenge of theoretical physics. While general relativity provides a description of gravity on macroscopic scales [Misner:1973prb], its prediction of spacetime singularities within black holes and at the cosmological origin marks the definitive boundary of its domain [Penrose:1964wq, Hawking:1970zqf]. The occurrence of infinite curvature indicates a breakdown of classical general relativity in the high-energy regime, making the inclusion of quantum effects essential for a consistent description. Understanding how such quantum corrections resolve the singularity is therefore a central task, and black holes provide a natural setting for investigating these issues. Within this context, the semiclassical approximation has yielded profound insights, most notably the discovery of black hole thermodynamics [Bekenstein:1973ur, Hawking:1975vcx]. The concepts of Bekenstein-Hawking entropy and Hawking temperature has shown a remarkable link between horizon area, surface gravity, and the laws of thermodynamics, suggesting a statistical-mechanical origin for spacetime geometry. Recent developments have further strengthened this perspective, highlighting the continuing relevance of black hole thermodynamics to the understanding of spacetime and gravitational dynamics [Strominger:1996sh, Padmanabhan:2009vy, Ashtekar:1997yu, Jacobson:2015hqa]. However, this paradigm relies on a singular background while simultaneously predicting the black hole’s complete evaporation, a process that would inevitably involve the singular region and exacerbate the information loss paradox [Giddings:1995gd, Unruh:2017uaw]. Therefore, a framework where the singularity is resolved from the beginning is necessary for a self-consistent description of the black hole life cycle. This necessity has given rise to the study of regular black holes, which serve as phenomenological models to explore the physical consequences of singularity resolution. An additional motivation comes from the fate of inner horizons. Strong cosmic censorship links physical predictability to the instability of Cauchy horizons, and recent analyses show that even in regular or quantum-corrected black holes these inner horizons typically undergo exponential mass amplification under perturbations [Cardoso:2017soq, Bonanno:2020fgp, Carballo-Rubio:2021bpr]. Moreover, mass-inflation–type instabilities can arise even in dynamical spacetimes without a strict Cauchy horizon, provided an inner trapping horizon evolves slowly [Carballo-Rubio:2024dca]. These results reinforce the view that long-lived, nonsingular black-hole spacetimes require a controlled modification of the interior structure These effective geometries replace the central singularity with a regular, non-singular core [Bardeen68, Hayward:2005gi]. Recent studies have investigated the dynamical formation of regular black holes including thin-shell and Oppenheimer–Snyder collapse as well as fully dynamical scenarios [Bueno:2024dgm, Bueno:2024eig, Bueno:2025gjg]. This regularity involves the introduction of matter that violates the Null Energy Condition (NEC). [Visser:1995cc, Visser:1989kh, Morris:1988tu]. Regular rotating black holes have also been investigated as part of this broader effort to construct nonsingular spacetimes [Bambi:2013ufa, Azreg-Ainou:2014pra, Torres:2022twv]. In particular, the work of Franzin et al. has demonstrated certain rotating regular black hole solutions whose inner horizon has zero surface gravity and is, therefore, stable against mass inflation [Franzin:2022wai, Ghosh:2022gka]. Among other wide proposals of regular spacetime models [Frolov:2016pav, Frolov:2017dwy, Bardeen:2018frm, Carballo-Rubio:2018pmi, Cano:2018aod], the Simpson-Visser black bounce model [Simpson:2018tsi, Lobo:2020ffi] offers a particularly compelling framework due to its elegant and minimalist construction. Governed by a single regularization parameter $a$, which introduces a characteristic length scale, it provides an analytic, continuous interpolation between the Schwarzschild black hole $(a\rightarrow 0)$, a regular one-way wormhole known as a “black-bounce,” and a two-way traversable wormhole. The black-bounce case is particularly interesting, as it represents an object that is asymptotically indistinguishable from a classical black hole yet possesses a fundamentally different internal structure, replacing the singularity with a smooth bounce to another asymptotic region. The central idea motivating our present work is that a geometric alteration as fundamental as singularity resolution must manifest as a distinct thermodynamic signature. Its modified structure must be reflected in its stability and response to quantum fluctuations. A number of previous studies have investigated the thermodynamics of regular black holes, most notably within Bardeen and Hayward models [Tharanath:2014naa, Maluf:2018lyu, Wu:2024gqi, PhysRevD.99.024015]. The Simpson–Visser ‘black-bounce’ metric forms a broader and qualitatively different class of regular black holes, interpolating smoothly between the Schwarzschild solution and a Morris–Thorne wormhole. As emphasized in the original work by Simpson and Visser [Simpson:2018tsi], this geometry generalizes and broadens the space of regular black hole solutions beyond the models usually considered. This difference in geometric structure leads to distinct thermodynamic behaviour, including the emergence of a Davies-type phase transition and a modified entropy structure, which motivates a dedicated analysis separate from previous studies. Our investigation therefore, proceeds along two complementary lines of inquiry. First, we conduct the analysis of the black hole’s thermodynamic stability. We go beyond the calculation of temperature and entropy to probe the system for critical behavior by examining the heat capacity and free energy as functions of the regularization parameter $a$. Our analysis reveals the existence of a critical point where the heat capacity diverges, signaling a second-order phase transition. This result provides concrete evidence that the regularization scale acts as a control parameter that can drive the system between phases of stability and instability. Such phase transitions where the heat capacity tends to change sign at a turning point are also in other black hole models [Jing:2008an, Croney:2025cue] Second, we advance beyond the semiclassical area law to quantify the leading-order quantum corrections to the black hole entropy. While the Bekenstein-Hawking entropy is a semiclassical result, the logarithmic corrections, which encode information about the quantum state counting and fluctuations of the background geometry itself, are a crucial part of the full quantum picture. Employing the well-established Hamilton-Jacobi tunneling method [Parikh:1999mf, Parikh:2004ih, Banerjee:2008cf, Majhi:2009pr], we explicitly derive these corrections for the Simpson-Visser spacetime. This calculation provides a refined statistical basis for the entropy of a regular black hole and serves as a crucial step towards understanding the quantum structure of non-singular spacetimes. Taken together, this dual analysis of classical thermodynamic stability and leading-order quantum corrections provides a complete thermodynamic portraits of a regular black hole. It demonstrates quantitatively how the removal of the singularity governs the object’s phase structure and modifies its quantum entropy, offering critical insights into the nature of black hole remnants and the resolution of the information paradox.
@@ III Thermodynamic Stability and Phase Transition
The thermodynamic response of a black hole, encapsulated by its heat capacity, provides deep insights into its stability and its ability to achieve equilibrium with a thermal environment. In a canonical ensemble, a positive heat capacity is a necessary condition for thermodynamic stability, whereas a negative capacity signals an instability. [Davies:1978]. The Schwarzschild black hole, with its negative heat capacity, becomes hotter as it radiates energy. Black hole phase transitions, including Van der Waals–like behavior and critical phenomena, have been widely analyzed in both singular and regular spacetimes [Witten:1998zw, Kastor:2009wy, Ali:2019rjn, Ruppeiner:2018pgn, Kubiznak:2012wp]. In this section, we investigate how the regularization parameter $a$ of the Simpson–Visser model fundamentally alters this behavior, revealing the existence of a critical point where the black hole undergoes a second-order phase transition from an unstable to a stable thermodynamic phase.
@@ III.2 Stability Analysis and Second-Order Phase Transition
This critical point lies within the allowed range for a regular black hole, $0<a_{\text{crit}}<2m$. The divergence of the heat capacity is the characteristic of a second-order phase transition [Davies:1977]. This transition separates two distinct thermodynamic phases, whose stability is dictated by the sign of the heat capacity.
@@ III.2.2 The Stable Phase ($\sqrt{2}m<a<2m$)
When the regularization parameter exceeds the critical value, the denominator $2m^{2}-a^{2}$ becomes negative. The heat capacity becomes positive ($C>0$). Consequently, black holes in this phase are thermodynamically stable. They can achieve stable thermal equilibrium with a surrounding heat bath, absorbing and emitting radiation to maintain a constant temperature. This behavior is a radical departure from the classical black hole picture and is a direct physical consequence of the modified spacetime geometry near the regular core.
The transition at $a=a_{\text{crit}}$ therefore represents a fundamental shift in the black hole’s physical nature. The regularization parameter $a$ is not merely a geometric cutoff but acts as a thermodynamic control parameter that drives the system across a critical point, transforming it from an unstable, perpetually evaporating object into a stable one. This thermodynamic behavior is illustrated in Fig:(2)
@@
Figure 3: The free energy indicates the thermodynamically favored state of the system. The plot reveals that the free energy is a monotonically increasing function of $a$, achieving its global minimum of $F=m/2$ at $a=0$. This implies that the singular Schwarzschild configuration is the thermodynamically preferred state among all regular black hole solutions of the same mass. While configurations with $a>\sqrt{2}m$ are locally stable (possessing a positive heat capacity), their higher free energy may make them globally disfavored. The vertical dashed line marks the critical point $a_{c}=\sqrt{2}m$ of the phase transition for reference.
@@ V Discussion and Conclusion
Our first major result was the discovery of a order phase transition at a critical regularization scale, $a_{\text{crit}}=\sqrt{2}m$. This transition separates an unstable, Schwarzschild-like thermodynamic phase ($a<a_{\text{crit}}$) from a thermodynamic stable phase ($a>a_{\text{crit}}$). This result implies that the regularization of the singularity fundamentally restructures the black hole’s thermodynamic landscape, permitting it to achieve stable thermal equilibrium with a surrounding environment, heat bath, a possibility forbidden for its singular counterpart. However, our analysis of the free energy added a crucial layer of physical nuance. We found that the free energy is minimized at $a=0$, indicating that the singular Schwarzschild state is the globally preferred thermodynamic configuration. This creates a compelling physical picture: the stable regular black holes can be interpreted as they are locally stable against small thermal fluctuations.
@@
The second result of our investigation was the derivation of the quantum-corrected entropy via the Hamilton-Jacobi tunneling formalism. Here, our analysis yielded several insights. Perhaps the most subtle but profound outcome was that the leading-order entropy, when derived through employing Hamilton-Jacobi approach, explicitly depends on the regularization parameter $a$. These results show that the thermodynamics is sensitive to the presence of the regular core from the very beginning, contradicting the expectation that the core only modifies subleading corrections. Such dependence is concealed if one applies the Bekenstein–Hawking area law directly at the horizon $r_{+}$. Furthermore, the higher-order correction term, scaling as $1/a^{2}$. Its divergence for small $a$ signals that the perturbative quantum expansion becomes incomplete precisely when the regular core is very small and the spacetime approaches its singular limit. This suggests that the regularization parameter $a$ functions not only as a geometric regulator but also as a regulator for the validity of the effective field theory description of the quantum entropy.
@@
Synthesizing these two threads, the thermodynamic phase structure and the corrected entropy, provides a powerful, self-consistent narrative for the end-stage of black hole evaporation. A large black hole initially in the unstable phase ($a\ll m$) will radiate and shrink. As its mass decreases, it eventually crosses the critical point into the stable phase, where its evaporation would dramatically slow. The process ceases when the black hole reaches the zero-temperature extremal state at $a=2m$, leaving a stable remnant. Our work provides the full statistical-mechanical description of this remnant. Its entropy, $S_{\text{remnant}}\propto\beta_{1}\ln(a)$, is a non-zero, purely logarithmic quantity determined entirely by the quantum gravity scale $a$ and the leading quantum correction coefficient $\beta_{1}$. This provides a concrete mechanism for resolving the information loss paradox: information is not lost to a singularity but is preserved in the quantum microstates of a stable, non-singular, Planck-scale object.

### v2

@@ I Introduction
The unresolved tension between general relativity and quantum mechanics stands as the central challenge of theoretical physics. While general relativity provides a description of gravity on macroscopic scales [Misner:1973prb], its prediction of spacetime singularities within black holes and at the cosmological origin marks the definitive boundary of its domain [Penrose:1964wq, Hawking:1970zqf]. The occurrence of infinite curvature indicates a breakdown of classical general relativity in the high-energy regime, making the inclusion of quantum effects essential for a consistent description. Understanding how such quantum corrections resolve the singularity is therefore a central task, and black holes provide a natural setting for investigating these issues. Within this context, the semiclassical approximation has yielded profound insights, most notably the discovery of black hole thermodynamics [Bekenstein:1973ur, Hawking:1975vcx]. The concepts of Bekenstein-Hawking entropy and Hawking temperature has shown a remarkable link between horizon area, surface gravity, and the laws of thermodynamics, suggesting a statistical-mechanical origin for spacetime geometry. Recent developments have further strengthened this perspective, highlighting the continuing relevance of black hole thermodynamics to the understanding of spacetime and gravitational dynamics [Strominger:1996sh, Padmanabhan:2009vy, Ashtekar:1997yu, Jacobson:2015hqa]. However, this paradigm relies on a singular background while simultaneously predicting the black hole’s complete evaporation, a process that would inevitably involve the singular region and exacerbate the information loss paradox [Giddings:1995gd, Unruh:2017uaw]. Therefore, a framework where the singularity is resolved from the beginning is necessary for a self-consistent description of the black hole life cycle. This necessity has given rise to the study of regular black holes, which serve as phenomenological models to explore the physical consequences of singularity resolution. An additional motivation comes from the fate of inner horizons. Strong cosmic censorship links physical predictability to the instability of Cauchy horizons, and recent analyses show that even in regular or quantum-corrected black holes these inner horizons typically undergo exponential mass amplification under perturbations [Cardoso:2017soq, Bonanno:2020fgp, Carballo-Rubio:2021bpr]. Moreover, mass-inflation–type instabilities can arise even in dynamical spacetimes without a strict Cauchy horizon, provided an inner trapping horizon evolves slowly [Carballo-Rubio:2024dca]. These results reinforce the view that long-lived, nonsingular black-hole spacetimes require a controlled modification of the interior structure. These effective geometries replace the central singularity with a regular, non-singular core [Bardeen68, Hayward:2005gi]. Recent studies have investigated the dynamical formation of regular black holes including thin-shell and Oppenheimer–Snyder collapse as well as fully dynamical scenarios [Bueno:2024dgm, Bueno:2024eig, Bueno:2025gjg]. This regularity involves the introduction of matter that violates the Null Energy Condition (NEC). [Visser:1995cc, Visser:1989kh, Morris:1988tu]. Regular rotating black holes have also been investigated as part of this broader effort to construct nonsingular spacetimes [Bambi:2013ufa, Azreg-Ainou:2014pra, Torres:2022twv]. In particular, the work of Franzin et al. has demonstrated certain rotating regular black hole solutions whose inner horizon has zero surface gravity and is, therefore, stable against mass inflation [Franzin:2022wai, Ghosh:2022gka]. Among other wide proposals of regular spacetime models [Frolov:2016pav, Frolov:2017dwy, Bardeen:2018frm, Carballo-Rubio:2018pmi, Cano:2018aod], the Simpson-Visser black bounce model [Simpson:2018tsi, Lobo:2020ffi] offers a particularly compelling framework due to its elegant and minimalist construction. Governed by a single regularization parameter $a$, which introduces a characteristic length scale, it provides an analytic, continuous interpolation between the Schwarzschild black hole $(a\rightarrow 0)$, a regular one-way wormhole known as a “black-bounce,” and a two-way traversable wormhole. The black-bounce case is particularly interesting, as it represents an object that is asymptotically indistinguishable from a classical black hole yet possesses a fundamentally different internal structure, replacing the singularity with a smooth bounce to another asymptotic region. The central idea motivating our present work is that a geometric alteration as fundamental as singularity resolution must manifest as a distinct thermodynamic signature. Its modified structure must be reflected in its stability and response to quantum fluctuations. A number of previous studies have investigated the thermodynamics of regular black holes, most notably within Bardeen and Hayward models [Tharanath:2014naa, Maluf:2018lyu, Wu:2024gqi, PhysRevD.99.024015]. The Simpson–Visser ‘black-bounce’ metric forms a broader and qualitatively different class of regular black holes, interpolating smoothly between the Schwarzschild solution and a Morris–Thorne wormhole. As emphasized in the original work by Simpson and Visser [Simpson:2018tsi], this geometry generalizes and broadens the space of regular black hole solutions beyond the models usually considered. This difference in geometric structure leads to distinct thermodynamic behaviour, including the emergence of a Davies-type phase transition and a modified entropy structure, which motivates a dedicated analysis separate from previous studies. Our investigation therefore, proceeds along two complementary lines of inquiry. First, we conduct the analysis of the black hole’s thermodynamic stability. We go beyond the calculation of temperature and entropy to probe the system for critical behavior by examining the heat capacity and free energy as functions of the regularization parameter $a$. Our analysis reveals the existence of a critical point where the heat capacity diverges, signaling a thermodynamic stability transition marked by a Davies point. This result provides concrete evidence that the regularization scale acts as a control parameter that can drive the system between phases of stability and instability. Such phase transitions where the heat capacity tends to change sign at a turning point are also in other black hole models [Jing:2008an, Croney:2025cue] Second, we advance beyond the semiclassical area law to quantify the leading-order quantum corrections to the black hole entropy. While the Bekenstein-Hawking entropy is a semiclassical result, the logarithmic corrections, which encode information about the quantum state counting and fluctuations of the background geometry itself, are a crucial part of the full quantum picture. Employing the well-established Hamilton-Jacobi tunneling method [Parikh:1999mf, Parikh:2004ih, Banerjee:2008cf, Majhi:2009pr], we explicitly derive these corrections for the Simpson-Visser spacetime. This calculation provides a refined statistical basis for the entropy of a regular black hole and serves as a crucial step towards understanding the quantum structure of non-singular spacetimes. Taken together, this dual analysis of classical thermodynamic stability and leading-order quantum corrections provides a complete thermodynamic portrait of a regular black hole. It demonstrates quantitatively how the removal of the singularity governs the object’s phase structure and modifies its quantum entropy, offering critical insights into the nature of black hole remnants and the resolution of the information paradox.
@@ III Thermodynamic Stability and Phase Transition
The thermodynamic response of a black hole, encapsulated by its heat capacity, provides deep insights into its stability and its ability to achieve equilibrium with a thermal environment. In a canonical ensemble, a positive heat capacity is a necessary condition for thermodynamic stability, whereas a negative capacity signals an instability. [Davies:1978]. The Schwarzschild black hole, with its negative heat capacity, becomes hotter as it radiates energy. Black hole phase transitions, including Van der Waals–like behavior and critical phenomena, have been widely analyzed in both singular and regular spacetimes [Witten:1998zw, Kastor:2009wy, Ali:2019rjn, Ruppeiner:2018pgn, Kubiznak:2012wp]. In this section, we investigate how the regularization parameter $a$ of the Simpson–Visser model fundamentally alters this behavior, revealing that this regular spacetime exhibits a critical instability, marked by a Davies point where the system undergoes a thermodynamic stability transition.”.
@@ III.2 Stability Analysis and Second-Order Phase Transition
This critical point lies within the allowed range for a regular black hole, $0<a_{\text{crit}}<2m$. The divergence of the heat capacity is the characteristic of a second-order phase transition [Davies:1977, Sokolowski:1980uva]. This transition separates two distinct thermodynamic phases, whose stability is dictated by the sign of the heat capacity.
@@ III.2.2 The Stable Phase ($\sqrt{2}m<a<2m$)
When the regularization parameter exceeds the critical value, the denominator $2m^{2}-a^{2}$ becomes negative. The heat capacity becomes positive ($C>0$). Consequently, black holes in this phase are thermodynamically stable. They can achieve stable thermal equilibrium with a surrounding heat bath, absorbing and emitting radiation to maintain a constant temperature. This behavior is a departure from the classical black hole picture and is a direct physical consequence of the modified spacetime geometry near the regular core.
The transition at $a=a_{\text{crit}}$ therefore represents a shift in the black hole’s physical nature. The regularization parameter $a$ is not merely a geometric cutoff but acts as a parameter that drives the system across a critical point, transforming it from an unstable, perpetually evaporating object into a stable one. This thermodynamic behavior is illustrated in Fig:(2)
@@
Figure 3: The free energy indicates the thermodynamically favored state of the system. The plot reveals that the free energy is an increasing function of $a$, achieving its global minimum of $F=m/2$ at $a=0$. This implies that the singular Schwarzschild configuration is the thermodynamically preferred state among all regular black hole solutions of the same mass. The physical thermodynamic evolution is instead governed by mass loss due to Hawking evaporation at fixed $a$. As the mass $m$ decreases, the system crosses a critical point at $m=a/\sqrt{2}$, beyond which the heat capacity becomes positive. This signals a change in the thermal response. The vertical dashed line marks the critical point $a_{c}=\sqrt{2}m$ of the phase transition for reference.
@@ V Discussion and Conclusion
Our first major result was the discovery of a Davies Point transition at a critical regularization scale, $a_{\text{crit}}=\sqrt{2}m$. This transition separates an unstable, Schwarzschild-like thermodynamic phase ($a<a_{\text{crit}}$) from a thermodynamic stable phase ($a>a_{\text{crit}}$). This result implies that the regularization of the singularity fundamentally restructures the black hole’s thermodynamic landscape, permitting it to achieve stable thermal equilibrium with a surrounding environment, heat bath, a possibility forbidden for its singular counterpart. However, our analysis of the free energy added a crucial layer of physical nuance. We found that the free energy is minimized at $a=0$, indicating that the singular Schwarzschild state is the preferred thermodynamic configuration. However, this minima is physically inaccessible if $a$ represents a fundamental length scale of the spacetime. In this scenario, the black hole cannot decay to the singular state. Instead, its thermodynamic evolution is driven by mass loss due to evaporation. As the mass m decreases for a fixed a, the system is driven towards the thermodynamically stable regime $(a>\sqrt{2}m)$, where the heat capacity becomes positive.
@@
The second result of our investigation was the derivation of the quantum-corrected entropy via the Hamilton-Jacobi tunneling formalism. Here, our analysis yielded several insights. Perhaps the most subtle but interesting outcome was that the leading-order entropy, when derived through employing Hamilton-Jacobi approach, explicitly depends on the regularization parameter $a$. These results show that the thermodynamics is sensitive to the presence of the regular core from the very beginning, contradicting the expectation that the core only modifies subleading corrections. Such dependence is concealed if one applies the Bekenstein–Hawking area law directly at the horizon $r_{+}$. Furthermore, the higher-order correction term, scaling as $1/a^{2}$. Its divergence for small $a$ signals that the perturbative quantum expansion becomes incomplete precisely when the regular core is very small and the spacetime approaches its singular limit. This suggests that the regularization parameter $a$ functions not only as a geometric regulator but also as a regulator for the validity of the effective field theory description of the quantum entropy.
@@
Synthesizing these two threads, the thermodynamic phase structure and the corrected entropy, provides a powerful, self-consistent narrative for the end-stage of black hole evaporation. A large black hole initially in the unstable phase ($a\ll m$) will radiate and shrink. As its mass decreases, it eventually crosses the critical point into the stable phase, where its evaporation would dramatically slow. The process ceases when the black hole reaches the zero-temperature extremal state at $a=2m$, leaving a stable remnant. Our work provides the statistical-mechanical description of this remnant. Its entropy, $S_{\text{remnant}}\propto\beta_{1}\ln(a)$, is a non-zero, purely logarithmic quantity determined entirely by the quantum gravity scale $a$ and the leading quantum correction coefficient $\beta_{1}$.